OpenStudy (lovely_lovely):

State whether each extreme in the data set is an outlier. A. The lower extreme is an outlier, but the upper extreme is not. B. The upper extreme is an outlier, but the lower extreme is not. C. Both extremes are outliers. D. Neither extreme is an outlier

OpenStudy (lovely_lovely):

State whether each extreme in the data set is an outlier. A. The lower extreme is an outlier, but the upper extreme is not. B. The upper extreme is an outlier, but the lower extreme is not. C. Both extremes are outliers. D. Neither extreme is an outlier

OpenStudy (lovely_lovely):

need help

OpenStudy (orange_):

read this...it will help http://www.purplemath.com/modules/boxwhisk3.htm

OpenStudy (orange_):

Q1 = 14 Q3 = 20 IQR = Q3 - Q1 = 20 - 14 = 6 Outliers will be below 14 - 1.5 * 6 = 14 - 9 = 5 or above 20 + 1.5 * 6 = 20 + 9 = 29. So your outliers are : 2 and 32....both of them (C)

OpenStudy (orange_):

Could I have a medal?

OpenStudy (orange_):

Just click "Best Response"

OpenStudy (orange_):

@lovely_lovely